OpenStudy (anonymous):

Find the exact number of ways to arrange 3 red marbles and 11 green marbles

OpenStudy (uri):

Is this the whole question?

OpenStudy (anonymous):

Yes. It's a statistics problem

OpenStudy (uri):

3! =3*2*1= 11! =11*10*9*8*7*6*5*4*3*2*1=